# Remote Auditing: Design-Based Tests of Randomization, Selection, and Missingness with Broadly Accessible Satellite Imagery

### v1

@@ Abstract
Randomized controlled trials (RCTs) are the benchmark for causal inference, yet field implementation can deviate. We here present a remote audit—a design-based, preregistrable diagnostic that uses only pre-treatment satellite imagery to test whether assignment is independent of local conditions. The conditional randomization test of the remote audit evaluates whether treatment assignment is more predictable from pre-treatment satellite features than expected under the experiment’s registered mechanism, providing a finite-sample valid, design-based diagnostic that requires no parametric assumptions. The procedure is finite-sample valid, honors blocks and clusters, and controls multiplicity across image models and resolutions via a max-statistic. We illustrate with two RCTs: Uganda’s Youth Opportunities Program, where the audit corroborates randomization and flags selection and missing-data risks; and a school-based trial in Bangladesh, where assignment is highly predictable from pre-treatment features relative to the stated design, consistent with independent concerns about irregularities. Remote audits complement balance tests, lower early-stage costs, and enable rapid design checks when baseline surveys are expensive or infeasible.
@@
Keywords: Randomization integrity; Satellite imagery; Conditional randomization test; Field experiments
Introduction
“Nullius in verba.”—Motto of the Royal Society (from Horace, Epistles I.i)
@@ Introduction
Randomized experiments have transformed empirical social science by offering credible causal leverage in hard-to-study environments (Rubin, 2005). In practice, however, the path from a pre-registered randomization mechanism to realized treatment assignment is sometimes fraught. Geography, logistics, bureaucratic discretion, and political pressures can all perturb assignment away from the intended design (Glennerster and Takavarasha, 2013; Olken, 2015). Even modest deviations can matter for inference, particularly when assignment correlates with contextual features that also shape outcomes (Battisti, 2017). Traditional safeguards such as centralized randomization draws, sealed lists, enumerator training, together with ex post diagnostics on covariate balance, manipulation checks, are indispensable (Morgan and Rubin, 2012; Bruhn and McKenzie, 2009). Yet they can be expensive, delayed, or underpowered in the very settings where field experiments are most valuable: low-resource environments, multi-jurisdiction programs, or government deployments in which baseline surveys are difficult or expensive to field at scale (Bouguen et al., 2019).
@@
We here propose a complementary tool: a remote audit of randomization integrity that relies solely on pre-treatment satellite imagery. The idea is simple. If treatment assignment is truly as-if random under the registered mechanism, then it should not be predictable from covariates extracted from images collected prior to randomization. Conversely, if implementers implicitly targeted more accessible, wealthier, less conflict-prone, or otherwise distinctive places—attributes that often leave visual traces at even moderate resolutions—then a predictive signal should be detectable in the imagery. Because satellite archives are global, historical, and increasingly easy to process, this diagnostic can be run early and cheaply, in parallel with—or well in advance of—baseline data collection.
@@
Our approach operationalizes this idea as a conditional randomization test (CRT) (Candes et al., 2018; Hennessy et al., 2016) tailored to field experiments. We (i) extract features from strictly pre-treatment images (e.g., Landsat, Sentinel) using interpretable indices (e.g., nightlight) or off-the-shelf backbones (e.g., CLIP-like encoders, ViT, Swin), (ii) train a predictive model of treatment using only these pre-treatment embeddings, summarize fit with an out-of-sample log-likelihood improvement statistic, and (iii) compare the observed statistic to its finite-sample reference distribution obtained by resampling from the known randomization scheme (honoring blocks, clusters, and treatment fractions). We further provide a max-statistic procedure (Westfall and Young, 1993) to adjust inference across multiple image models, resolutions, and patch sizes, and we discuss simple alternatives like Bonferroni or BH-style control (Thissen, Steinberg and Kuang, 2002).
@@
The remote audit is design-based: validity does not hinge on outcome models or parametric assumptions, only on resampling from the registered assignment mechanism. It uses no post-treatment variables, thereby avoiding “bad control” pitfalls (Angrist and Pischke, 2009; Pearl, 2009; Cinelli and Hazlett, 2020), and because we never condition on image features in the outcome model, we avoid mediator/collider concerns that arise if images are measured after treatment. The audit complements standard balance tests: whereas balance checks examine low-dimensional covariates (often unavailable ex ante), the audit leverages high-dimensional, pre-existing visual context that is ubiquitous and pre-treatment.
@@
Our contribution is fourfold. First, we formalize a preregistrable CRT for randomization integrity explicitly adapted to geospatial experiments and to off-the-shelf computer vision embeddings. Second, we provide practical guidance on implementation: pre-treatment image selection, patching and scale, sample-splitting to prevent overfitting, multiple-testing control across models/resolutions, and transparent reporting. Third, we show empirically that remote audits can be informative where baseline covariates are unavailable, delayed, or incomplete—a common situation in development and governance experiments (Dreher and Lohmann, 2015; BenYishay, DiLorenzo and Dolan, 2022; Weisberg, 2009). Fourth, we release a free and coding-free audit online computer application that experimenters can use to perform audits at scale (URL: https://huggingface.co/spaces/cjerzak/ra).
In brief, remote audits will not detect all implementation problems. Many political and social processes are not visible from space, and clouds, revisit cycles, and spatial resolution impose constraints. Nonetheless, as a minimally invasive, design-based check, a remote audit can serve as an early warning system, flagging potential risks and guiding remedial steps before costly downstream investments.
@@ 1 A Conditional Randomization Test for Remote Audits
1.1 Design Setup
@@
Let $\Omega$ denote the registered randomization procedure (e.g., complete randomization at rate $\bar{a}$, or stratified randomization within blocks with fixed treatment counts). Consider experimental units $i=1,\ldots,n$ (e.g., villages, neighborhoods, clinics), each with geospatial coordinates $\ell_{i}\in\mathbb{R}^{2}$. Prior to any intervention, we extract a pre-treatment image patch of size $s>0$ centered on $\ell_{i}$ from a public image archive (e.g., Landsat), denoted $\mathbf{M}_{i}=f_{M}(\ell_{i},s)$. From $\mathbf{M}_{i}$, we compute a representation, $\boldsymbol{\phi}_{i}=f_{\boldsymbol{\phi}}(\mathbf{M}_{i})\in\mathbb{R}^{d}$ using either interpretable indices (e.g., vegetation or texture), task-agnostic image embedding models (e.g., CLIP-like encoders, ViT, Swin), or both. [^1] Embeddings are generated variables. When used as regressors for causal estimation they can require additional care (Battaglia et al., 2024). Our design-based test avoids such complications because we only use embeddings to form a test statistic for assignment, not to adjust outcome models. Let $A_{i}\in\{0,1\}$ denote treatment assignment and $Y_{i}$ the outcome.
@@
Under the experiment’s intended design, assignment is independent of all pre-treatment variables, including $\boldsymbol{\phi}_{i}$:
$A_{i}\;\perp\!\!\!\perp\;\boldsymbol{\phi}_{i}$
@@
conditional on the known randomization scheme $\Omega$ (e.g., complete randomization at rate $\bar{a}$, or stratified/block randomization within pre-defined strata). Informally, $\boldsymbol{\phi}_{i}$ should not help predict $A_{i}$ beyond the treatment fraction implied by $\Omega$. If $\boldsymbol{\phi}_{i}$ does predict $A_{i}$ in the observed data substantially better than it typically does under draws from $\Omega$, that is evidence that the realized assignment is atypical of the design—consistent with implementer discretion, operational constraints, or administrative errors aligning treatment with visual correlates of local conditions.
In contrast to covariate balance tests, which examine a low-dimensional set of pre-specified variables, remote audits leverage high-dimensional pre-treatment information available almost everywhere on Earth. [^2] Earth-observation missions provide global coverage and long archives, though revisit, cloud cover, and licensing constraints matter (Barnum, 2022; Tao et al., 2016; Townsend, 2021). Because $\mathbf{M}_{i}$ is measured before treatment on a daily basis for all geographical areas, we avoid post-treatment bias (Angrist and Pischke, 2009; Pearl, 2009); because we never condition on $\boldsymbol{\phi}_{i}$ in outcome analysis, we avoid introducing bad controls or colliders.
@@
Figure 1: Remote audit intuition. Is latent context $U_{i}$, proxied by imagery, correlated with treatment assignment? The dark, thicker arrow marks the edge whose (effective) presence would make assignment $A_{i}$ predictable from pre-treatment imagery via embeddings, i.e., the path $\boldsymbol{\phi}_{i}\!\to\!A_{i}$. Under the registered mechanism, $A_{i}\perp\!\!\!\perp\boldsymbol{\phi}_{i}$.
1.2 Test Statistic
To turn this intuition into a test, we need a single, learner-agnostic statistic that captures how much the pre-treatment image information predicts treatment assignment beyond what the registered mechanism $\Omega$ would yield by chance. The desiderata are simple: evaluate strictly out of sample to prevent overfitting; anchor the scale to $\Omega$ so that “no signal” maps asymptotically to zero; allow additivity across units and folds for transparent cross-fitting; and avoid dependence on any particular classifier. A natural choice satisfying these criteria is the improvement in the predictive log-likelihood of $A$ on held-out data relative to the baseline assignment probabilities implied by $\Omega$ (or, in stratified/clustered designs, the unit-level baselines $q_{i}$). We treat the embedding–learner pair as a black box that outputs predicted probabilities and summarize only their incremental information. We now define this statistic formally; the Appendix derives its finite-sample reference distribution under $\Omega$.
First, split the sample at random into train and test folds (or use $K$-fold cross-fitting). Using only the training fold, fit any predictive model of $A_{i}$ on $\boldsymbol{\phi}_{i}$ (e.g., ridge-penalized logistic regression, gradient-boosted trees, or a neural model). Denote the probability predictions on the test fold by $\hat{\pi}_{i}=\widehat{\Pr}(A_{i}=1\mid\boldsymbol{\phi}_{i})$. Define the out-of-sample log-likelihood:
$\mathcal{L}\;=\;\sum_{i\in\text{test}}\left(A_{i}\log\hat{\pi}_{i}+(1-A_{i})\log(1-\hat{\pi}_{i})\right).$
To facilitate interpretation, subtract the baseline log-likelihood under the marginal treatment probability $\bar{a}$ implied by $\Omega$:
$T\;=\;\mathcal{L}\;-\;\sum_{i\in\text{test}}\big(A_{i}\log\bar{a}+(1-A_{i})\log(1-\bar{a})\big).$ (1)
Under genuine randomization, we expect $T$ to approach zero (from below) as the learner predicts the baseline treatment assignment with more precision (so that $T\to\mathcal{L}-\mathcal{L}=0$). Substantially positive values indicate that $\boldsymbol{\phi}_{i}$ predicts $A_{i}$ better than chance.

### v2

@@ Abstract
Randomized controlled trials (RCTs) are the benchmark for causal inference, yet field implementation can drift from the registered design or, by chance, yield imbalances. We introduce a remote audit—a preregistrable, design‑based diagnostic that uses strictly pre‑treatment, publicly available satellite imagery to test whether assignment is independent of local conditions. The audit implements a conditional randomization test that asks whether treatment is more predictable from pre‑treatment features than under the registered mechanism, delivering a finite‑sample‑valid, nonparametric check that honors blocks and clusters and controls multiplicity across image models, resolutions, and patch sizes via a max‑statistic. The same preregistered procedure can be run before baseline data collection to guide implementation and, after assignments are realized, to audit the actual allocation. In two illustrations—Uganda’s Youth Opportunities Program (randomization corroborated) and a school‑based experiment in Bangladesh (assignment predictable relative to the design, consistent with independent concerns)—the audit can surface potential problems early, before costly scientific investments. We also provide descriptive diagnostics for selection into the study and for missingness. Because it is low‑cost and can be implemented rapidly in a unified way across diverse global administrative jurisdictions, the remote audit complements balance tests, strengthens preregistration, and enables rapid design checks when conventional data collection is slow, expensive, or infeasible.
@@
Keywords: Field experiments; Satellite imagery; Conditional randomization test
Word count: 6,057
@@
Randomized experiments have transformed empirical social science by offering credible causal leverage in hard-to-study environments (Rubin, 2005; Baldassarri and Abascal, 2017; Gerber and Green, 2017). In practice, however, the path from a pre-registered randomization mechanism to realized treatment assignment is sometimes fraught. Geography, logistics, bureaucratic discretion, political pressures, and bad randomization draws can all perturb assignment away from the intended design or away from covariate balance between treatment groups (Glennerster and Takavarasha, 2013; Olken, 2015). Even modest deviations can matter for inference, particularly when assignment correlates with contextual features that also shape outcomes (Battisti, 2017). Traditional safeguards such as centralized (re)randomization draws, sealed lists, enumerator training, together with ex-post diagnostics on covariate balance, manipulation checks, are indispensable (Morgan and Rubin, 2012; Bruhn and McKenzie, 2009). Yet they can be expensive, delayed, or underpowered in the very settings where field experiments are most valuable: low-resource environments, multi-jurisdiction programs, or government deployments in which baseline surveys are difficult or expensive to field at scale (Bouguen et al., 2019).
@@
We here propose a complementary tool for field experiments: a remote audit of randomization integrity that relies on pre-treatment satellite imagery. [^1] Our scope is field experiments implemented in real-world settings; we do not target laboratory studies or small within-organization trials (e.g., within-school classrooms). The idea is simple. Under the registered mechanism, treatment assignment should not be predictable from covariates extracted from images collected before randomization. If implementers implicitly targeted more accessible, wealthier, less conflict-prone, or otherwise distinctive places—attributes that often leave visual traces even at moderate resolution—then a predictive signal should be detectable in the imagery.
@@
Our approach operationalizes these ideas as a conditional randomization test (CRT) (Candes et al., 2018; Hennessy et al., 2016) tailored to field experiments. We (i) extract features from strictly pre-treatment images (e.g., Landsat, Sentinel) using interpretable indices (e.g., nightlight) or off-the-shelf backbones (e.g., CLIP-like encoders, ViT, Swin; Xiao et al. (2025)), (ii) train a predictive model of treatment using only these pre-treatment embeddings, summarize fit with an out-of-sample log-likelihood improvement statistic, and (iii) compare the observed statistic to its finite-sample reference distribution obtained by resampling from the known randomization scheme (honoring blocks, clusters, and treatment fractions). We further provide a max-statistic procedure (Westfall and Young, 1993) to adjust inference across multiple image models, resolutions, and patch sizes, and we discuss simple alternatives like Bonferroni or BH-style control (Thissen, Steinberg and Kuang, 2002).
@@
The remote audit of randomization quality is design-based: validity does not hinge on outcome models or parametric assumptions, only on resampling from the registered assignment mechanism. It uses no post‑treatment variables, mitigating “bad control” risks (Angrist and Pischke, 2009; Pearl, 2009); and because we never condition on image features in the outcome model, the audit sidesteps adjustment‑based mediator and collider concerns. Pre‑treatment measurement helps but is not, by itself, sufficient for universally valid adjustment—see Cinelli and Hazlett (2020)—which is precisely why we use imagery only to form a design‑based test of assignment. The audit complements standard balance tests: whereas balance checks examine low-dimensional covariates (often unavailable ex ante), the audit leverages high-dimensional, pre-existing visual context that is ubiquitous and pre-treatment.
The remote audit can be used both (i) ex ante (before or during field mobilization) to probe fidelity of implementation to stated design and guide remedial steps if issues are found, and (ii) ex post to evaluate the quality of a realized assignment vector. In this note, we focus on the pre‑treatment, randomization audit. We later describe two descriptive extensions—selection and missingness diagnostics—that can be run either ex ante or ex post, but are not design-valid tests in the same sense as for randomization. Operationally, these extensions retarget the same predictability exercise, asking whether pre‑treatment imagery predicts these labels better than a baseline. Because there is typically no registered mechanism for missingness or selection, we treat the resulting evidence as descriptive diagnostics rather than design‑valid tests, but the data, folds, and estimation machinery are identical to the randomization audit.
@@
Our contribution is threefold. First, we formalize a preregistrable conditional randomization test for randomization integrity explicitly adapted to experiments that leverage freely and globally available satellite imagery. Second, we provide a practical workflow—pre‑treatment image selection, patching and scale, out‑of‑sample evaluation, and multiplicity control—and release a no‑code application that implements the audit at scale. [^2] URL: https://audit.planetarycausalinference.org Third, we show empirically that remote audits are informative when baseline covariates are unavailable or delayed, illustrating randomization, selection, and missingness diagnostics in two prominent field experiments (Dreher and Lohmann, 2015; BenYishay, DiLorenzo and Dolan, 2022; Weisberg, 2009).
Although remote audits extend the scope of checks and balances to increase the quality of field experiments, they will not detect all implementation problems. Many political and social processes are not visible from space, and clouds, revisit cycles, and spatial resolution impose constraints. Nonetheless, as a minimally invasive, design-based check, a remote audit can serve as an early warning system, bringing to light potential risks and guiding remedial steps before costly and slow downstream scientific investments.
1 The Remote Audit in Context: Related Work
Researchers already use satellite imagery to measure outcomes, build covariates, and study heterogeneity (Jean et al., 2016; Yeh et al., 2020; Jerzak, Johansson and Daoud, 2023; Torres and Pugh, 2022). Imagery has also been used in a model-based approach to mitigate confounding when pre-treatment signals proxy latent conditions (Sanford, 2021; Burke et al., 2021). Our contribution is orthogonal to these uses. We deploy imagery in a design-based capacity: a conditional randomization test that asks whether realized treatment assignment is independent of satellite-derived features under the registered mechanism. Because the audit never adjusts outcomes, it avoids “bad control” pitfalls and mediator/collider concerns that attend post-treatment measurement in model-based pipelines (Angrist and Pischke, 2009; Pearl, 2009; Cinelli and Hazlett, 2020). Where model-based approaches typically require identification assumptions and sensitivity analysis (Rosenbaum and Rubin, 1983; Oster, 2019), our audit derives validity from the design itself: extremeness is judged against the experiment’s finite-sample reference generated by draws from the assignment mechanism. In this sense, the paper reframes what imagery is—not as a control set within an outcome model, but as a ubiquitous, pre-existing source of design-stage information capable of certifying whether an assignment behaves as if it is random based on observables.
@@ 2 A Conditional Randomization Test for Remote Audits
2.1 Design & Data: Units, Imagery, Representations
@@
Let $\Omega$ denote the registered randomization procedure (e.g., complete randomization at rate $\bar{a}$, or stratified randomization within blocks with fixed treatment counts). Consider experimental units $i=1,\ldots,n$ (e.g., villages, neighborhoods, clinics), each with geospatial coordinates $\ell_{i}\in\mathbb{R}^{2}$. Let $A_{i}\in\{0,1\}$ denote treatment assignment and $Y_{i}$ the outcome. Prior to any intervention, we extract a pre-treatment image patch of size $s>0$ centered on $\ell_{i}$ from a public image archive (e.g., Landsat), denoted $\mathbf{M}_{i}=f_{M}(\ell_{i},s)$.
From $\mathbf{M}_{i}$, we compute a representation $\boldsymbol{\phi}_{i}=f_{\boldsymbol{\phi}}(\mathbf{M}_{i})\in\mathbb{R}^{d}$ using either interpretable indices (e.g., vegetation or texture) or off‑the‑shelf encoders (e.g., CLIP‑like, ViT, Swin), or both. Here, by representation we mean a fixed‑length numeric summary of an image that compresses visible patterns such as roads, settlement structure, vegetation, and roof materials into measurements usable downstream by standard predictive models. [^3] Embeddings are generated variables. When used as regressors for causal estimation, they can require additional care (Battaglia et al., 2024). Our design-based test avoids such complications because we only use embeddings to form a test statistic for assignment, not to adjust outcome models. We use a representation $\boldsymbol{\phi}_{i}=f_{\boldsymbol{\phi}}(\mathbf{M}_{i})$ rather than the raw image $\mathbf{M}_{i}$ to make evaluation computationally tractable; similar logic applies to both.
@@
Under the experiment’s intended design $\Omega$, assignment is independent of all pre-treatment variables, including any fixed representation of the image:
$A_{i}\;\perp\!\!\!\perp\;\mathbf{M}_{i}\quad\Rightarrow\quad A_{i}\;\perp\!\!\!\perp\;\boldsymbol{\phi}_{i},$
which is equivalent to the statement that, conditional on $\Omega$, $\boldsymbol{\phi}_{i}$ contains no information for predicting $A_{i}$ beyond the baseline treatment probabilities implied by the design. [^4] If blocking or stratification is used, independence should hold within clusters, not unconditionally.
For validity, we require only that $f_{\boldsymbol{\phi}}$ is fixed ex ante and that $\boldsymbol{\phi}_{i}$ is constructed strictly pre‑treatment from $\mathbf{M}_{i}$ captured before any mobilization. Under $\Omega$, $A_{i}\perp\!\!\!\perp\mathbf{M}_{i}$ implies $A_{i}\perp\!\!\!\perp\boldsymbol{\phi}_{i}$ for any fixed $f_{\boldsymbol{\phi}}$, so CRT pp‑values are finite‑sample valid regardless of how informative $\boldsymbol{\phi}_{i}$ is. For power, richer $\boldsymbol{\phi}_{i}$ (e.g., pretrained encoders plus interpretable indices) help detect departures when they exist, but discarding information can only reduce power, not invalidate the test.
@@
Informally, $\boldsymbol{\phi}_{i}$ should not help predict $A_{i}$ beyond the treatment fraction implied by $\Omega$. If $\boldsymbol{\phi}_{i}$ does predict $A_{i}$ in the observed data substantially better than it typically does under draws from $\Omega$, that is evidence that the realized assignment is atypical of the design—consistent with implementer discretion, operational constraints, or administrative errors aligning treatment with visual correlates of local conditions. In contrast to covariate balance tests of experimentor-collected features, which examine a low-dimensional set of pre-specified variables, remote audits can leverage high-dimensional pre-treatment information available almost everywhere on Earth. [^5] Earth-observation missions provide global coverage and long archives, though revisit, cloud cover, and licensing constraints matter (Barnum, 2022; Tao et al., 2016; Townsend, 2021).
Figure 1 encodes the estimand system at the level of raw pre‑treatment imagery: $\mathbf{M}_{i}$ (what satellites see before any intervention), assignment $A_{i}$, outcome $Y_{i}$, and latent context $U_{i}$ that shapes both what satellites see and potential outcomes. Under the registered mechanism there is no edge $\mathbf{M}_{i}\!\to\!A_{i}$, i.e., $A_{i}\perp\!\!\!\perp\mathbf{M}_{i}$. The audit asks whether the realized data behave as if an effective $\mathbf{M}_{i}\!\to\!A_{i}$ link were present. Substantively, such a link could arise if implementers followed a human map or administrative rule that favors, say, road-adjacent or wealthier-looking areas—patterns that $\mathbf{M}_{i}$ proxies even if those rules never referenced imagery explicitly. We subsequently use features or embeddings, $\boldsymbol{\phi}_{i}=f_{\boldsymbol{\phi}}(\mathbf{M}_{i})$, purely as a computational device to test for such a link.
@@
Figure 1: Remote audit intuition. Is latent context $U_{i}$, proxied by pre‑treatment imagery $\mathbf{M}_{i}$, correlated with treatment $A_{i}$? Under the registered mechanism, there is no $\mathbf{M}_{i}\!\to\!A_{i}$ edge; the CRT probes whether the realized assignment behaves as if such an edge were present.
2.2 Test Statistic: Out-of-Sample Likelihood Improvement
To turn this intuition into a test, we need a single, learner-agnostic statistic that captures how much the pre-treatment image information predicts treatment assignment beyond what the registered mechanism $\Omega$ would yield by chance. The desiderata are simple: evaluate strictly out of sample to prevent overfitting; anchor the scale to $\Omega$ so that “no signal” maps asymptotically to zero; allow additivity across units and folds for transparent cross-fitting; and avoid dependence on any particular classifier.
A natural choice satisfying these criteria is the improvement in the predictive log-likelihood of $A$ on held-out data relative to the baseline assignment probabilities implied by $\Omega$. Split the sample (or use $K$-fold cross-fitting). Fit a predictive model for $A_{i}$ using only $\boldsymbol{\phi}_{i}$ on the training fold; let $\hat{\pi}_{i}=\widehat{\Pr}(A_{i}=1\mid\boldsymbol{\phi}_{i})$ be its predictions on the test fold. Define:
$\mathcal{L}\;=\;\sum_{i\in\text{test}}\left(A_{i}\log\hat{\pi}_{i}+(1-A_{i})\log(1-\hat{\pi}_{i})\right),$
and let $\bar{a}$ be the marginal treatment rate under $\Omega$. Our test statistic is:
$(T=)\;\Delta\mathcal{L}\;\equiv\;\mathcal{L}\;-\;\sum_{i\in\text{test}}\big(A_{i}\log\bar{a}+(1-A_{i})\log(1-\bar{a})\big).$ (1)
Under genuine randomization, strictly out-of-sample fitting makes $\Delta\mathcal{L}$ concentrate near zero. In fact, when there is no signal, the baseline assignment probabilities implied by $\Omega$ (the global rate $\bar{a}$ are Bayes optimal: they maximize the expected held-out log-likelihood. Any learner that perturbs these baselines without real information will, by the concavity of $\log$, on average lose log-likelihood on test folds. Hence, $\mathbb{E}_{\Omega}[\Delta\mathcal{L}]\leq 0$ with equality only when $\hat{\pi}_{i}$ collapses to the baselines. Small negative values of $\Delta\mathcal{L}$ are therefore common in finite samples due to training noise and cross-fitting variability. Substantially positive values, by contrast, require genuine predictability from pre-treatment information; the Appendix details the finite-sample randomization reference under $\Omega$ used to quantify extremeness.